OpenStudy (anonymous):

What is the slope of the line 5x - 2y = -10?

Directrix (directrix):

Do you know how to write 5x - 2y = -10 in y = mx + b form?

OpenStudy (anonymous):

no

Directrix (directrix):

Take 5x - 2y = -10 and isolate the y term on the left. So, 5x - 5x - 2y = -10 -5x I added -5x to both sides. Your turn: Simplify: 5x - 5x - 2y = -10 -5x.

OpenStudy (anonymous):

x- 2y = -5x

Directrix (directrix):

(5x - 5x) - 2y = -10 -5x. Try again.

OpenStudy (anonymous):

-2y = -10 -5x

Directrix (directrix):

Good. Now, divide all terms of both sides of the equation by -2. -2y = -10 -5x I do it with reciprocals: (-1/2) * (-2y) = (-1/2)*(-10 - 5x) You may prefer a different but equivalent approach. Your turn: What is -2y = -10 -5x divided by -2?

OpenStudy (anonymous):

y = 5 -5x

Directrix (directrix):

All terms of -2y = -10 -5x have to be divided by -2. I agree with the y = 5 part but the -5x also have to be divided by -2. So, what is -5x divided by -2? Post and we'll compare notes.

OpenStudy (anonymous):

0? so x = 0

OpenStudy (anonymous):

no 2.5

OpenStudy (anonymous):

x=2.5

Directrix (directrix):

You're jumping ahead. We need to move step by step right now. -5x divided by -2 = (5/2)x So, now we have: y = 5 +(5/2) x which is y = (5/2) x + 5. Okay, Look at the slope-intercept form: y = mx + b and pick out what has to be to slope of y = (5/2) x + 5.

OpenStudy (anonymous):

5 over 2

Directrix (directrix):

Yes. That means that every time x moves up by 2, the corrresponding value of y moves up by 5.

OpenStudy (anonymous):

okay thanks

Directrix (directrix):

Glad to help.